OpenStudy (anonymous):

Solve for x. 6(x - 2) = 4

OpenStudy (anonymous):

that was fast

OpenStudy (anonymous):

x = 8/3 = 2.667

OpenStudy (anonymous):

x=1 x=1(1 / 3) x=2( 2 / 3) those are the answer choices

OpenStudy (anonymous):

i say 3 :/

OpenStudy (anonymous):

good luck

OpenStudy (anonymous):

3 or the 3rd choice?

OpenStudy (anonymous):

3rd choice

OpenStudy (anonymous):

alrighty, thanks